# with $f$-divergence Minimization

### v1

@@ Abstract
Preference optimization has made significant progress recently, with numerous methods developed to align language models with human preferences. This paper introduces $f$-divergence Preference Optimization ($f$-PO), a novel framework that generalizes and extends existing approaches. $f$-PO minimizes $f$-divergences between the optimized policy and the optimal policy, encompassing a broad family of alignment methods using various divergences. Our approach unifies previous algorithms like DPO and EXO, while offering new variants through different choices of $f$-divergences. We provide theoretical analysis of $f$-PO’s properties and conduct extensive experiments on state-of-the-art language models using benchmark datasets. Results demonstrate $f$-PO’s effectiveness across various tasks, achieving superior performance compared to existing methods on popular benchmarks such as AlpacaEval 2, Arena-Hard, and MT-Bench. Additionally, we present ablation studies exploring the impact of different $f$-divergences, offering insights into the trade-offs between regularization and performance in offline preference optimization. Our work contributes both practical algorithms and theoretical understanding to the field of language model alignment. Code is available at https://github.com/MinkaiXu/fPO.
@@ 2 Related Work
Aligning pretrained large language models with human preferences for high-quality responses is vital in natural language generation. RLHF (Christiano et al.,, 2017; Ouyang et al.,, 2022) has emerged as a common solution to address this problem by utilizing popular RL methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017a, ). However, this approach faces limitations due to training instability and the complexity introduced by its two-stage pipeline. DPO (Rafailov et al.,, 2023) overcomes this limitation by defining the preference loss as a function of the policy directly given the pairwise preference data, which directly optimizes the policy without an explicit surrogate reward and the RL phase. Followup works have extended this algorithm to utilize multiple ranked responses instead of pairwise preference data (Yuan et al., 2023b, ; Liu et al., 2024a, ; Song et al.,, 2024), and avoid dependence on the reference model, effectively merging the instruction tuning phase and preference optimization phase (Hong et al.,, 2024; Meng et al.,, 2024). GPO (Tang et al.,, 2024) unifies several DPO variants (Azar et al.,, 2023; Liu et al., 2024b, ) into a family of algorithms and provides an empirical analysis of the performance trade-off, but leave the theoretical analysis in a heuristic manner. Notably, Wang et al., (2023) also introduces $f$-divergence into the alignment problem, but the divergences are only applied to the regularization term of the RL formulation. Furthermore, all the above works are built on the reparameterization of reward function by policy likelihood, which cannot be guaranteed in practice (Ji et al.,, 2024). In contrast, our method unifies the algorithm based on the theoretical framework of statistical distribution divergence (Csiszár et al.,, 2004; Liese and Vajda,, 2006), which offers a principled to choose the specific algorithm based on divergence characteristics. EXO (Ji et al.,, 2024) proposes learning the policy by exactly optimizing the RLHF objective that minimizes the reverse KL divergence against the optimal policy, which is a special case of our proposed framework. In addition, numerous offline preference optimization works propose different training objectives to emphasize various behaviors (Yuan et al., 2023a, ; Zhao et al.,, 2023; Xu et al.,, 2024; Ethayarajh et al.,, 2024).
@@ 3.3 $f$-divergence
Divergence describes the difference between two probability distributions. A general family of divergences $f$-divergences (Csiszár et al.,, 2004; Liese and Vajda,, 2006), also known as the Ali-Silvey distances (Ali and Silvey,, 1966). For a convex function $f:\mathbb{R}^{+}\rightarrow\mathbb{R}$ that is lower-semicontinuous, strictly convex around $1$, and satisfies $f(1)=0$, given two distributions $p$ and $q$, the corresponding $f$-divergence for these two distributions is defined as:
@@ 4.2 Generalized Preference Optimization
Pair-wise Preference Data. The most common preference dataset typically consists of pair-wise completions for each prompt, i.e., setting $K=2$. Since there are two completions $y$ for each prompt $x$, we can denote them as “winning” sample $y_{w}$ and “losing” sample $y_{l}$. Then LABEL:eq:loss_fpo_general can be simplified as:
@@ 5.2 Benchmarking $f$-PO
In this subsection, we further demonstrate the strong performance of our $f$-PO on three popular benchmarks through four different settings.
@@
Evaluation metrics. We evaluate our models using three popular open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al.,, 2023; Dubois et al.,, 2024), MT-Bench (Zheng et al.,, 2023), and Arena-Hard 0.1 (Li et al.,, 2024). These benchmarks assess diverse conversational abilities across various queries. For AlpacaEval 2, we report both the raw win rates (WR) and length-controlled win rates (LC). MT-Bench scores are averaged using GPT-4 and GPT-4-Preview-1106 as judges. Arena-Hard results are reported as win rates against the baseline model.
@@
Results. Table 3 shows the primary evaluation outcomes of Mistral-7B and Llama-3-8B in both Base and Instruct configurations. While all methods effectively boost the performance over the SFT model, our proposed method, $f$-PO, achieves the highest score on 16 out of 20 metrics while being competitive on the rest, which highlights the efficacy and robustness of our technique. Quantitatively, $f$-PO achieves an average improvement of 1.4% over the strongest baseline SimPO on the AlpacaEval 2 length-controlled win rate metric. Interestingly, in the Llama-3-8B Base setup, we are able to obtain a remarkable enhancement of 4.8% (28.2% against 23.4%) in win rate compared with SimPO on Arena-Hard, further demonstrating the superiority of $f$-PO. The results consistently advocate $f$-PO as an effective and broadly applicable LLM alignment approach across a wide suite of models and tasks.
@@ 6 Conclusion
$f$-PO: Generalized Preference Optimization
with $f$-divergence Minimization
@@ A.2 Proof of Theorem 2
We first consider the following equation:
$\sum_{j=1}^{K}e^{g_{\theta}\left(x,y_{j}\right)}/K=\mathbb{E}_{\pi_{\textnormal{ref}}(y\mid x)}e^{g_{\theta}\left(x,y\right)}.$ (23)
When $K\to\infty$, the above equation holds because the LHS is the unbiased Monte Carlo estimation of the expectation. Note that $g_{\theta}(x,y)=\log\hat{\pi}_{\theta}(y\mid x)-\log\pi_{\mathrm{ref}}(y\mid x)=\log\frac{\hat{\pi}_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}$. Then we have the fact that:
@@ Appendix C Prompt Details
In this section, we provide the specific prompts used in our model evaluation, as described in Section 5.1 and Section 5.2.
@@ C.1 Dialogue Generation Evaluation Prompt
SECOND, on a new line, state only ”A” or ”B” to indicate which response is more helpful. Your response should use the format: Comparison: <one-sentence comparison and explanation>
More helpful: <”A” or ”B”>
@@ C.2 Summarization Evaluation Prompt
For assessing the quality of summaries, we employ the concise prompt structure adapted from Rafailov et al., (2023):
@@
SECOND, on a new line, state only ”A” or ”B” to indicate your choice. Your response should use the format: Comparison: <one-sentence comparison and explanation>
Preferred: <”A” or ”B”>

### v2

# Jiaqi Han, Mingjian Jiang, Yuxuan Song, Stefano Ermon, Minkai Xu∗†

@@ Abstract
Preference optimization has made significant progress recently, with numerous methods developed to align language models with human preferences. This paper introduces $f$-divergence Preference Optimization ($f$-PO), a novel framework that generalizes and extends existing approaches. $f$-PO minimizes $f$-divergences between the optimized policy and the optimal policy, encompassing a broad family of alignment methods using various divergences. Our approach unifies previous algorithms like DPO and EXO, while offering new variants through different choices of $f$-divergences. We provide theoretical analysis of $f$-PO’s properties and conduct extensive experiments on state-of-the-art language models using benchmark datasets. Results demonstrate $f$-PO’s effectiveness across various tasks, achieving superior performance compared to existing methods on popular benchmarks such as AlpacaEval 2, Arena-Hard, MT-Bench, and Open LLM Leaderboard v2. Additionally, we present ablation studies exploring the impact of different $f$-divergences, offering insights into the trade-offs between regularization and performance in offline preference optimization. Our work contributes both practical algorithms and theoretical understanding to the field of language model alignment. Code is available at https://github.com/MinkaiXu/fPO.
@@ 2 Related Work
Aligning pretrained large language models with human preferences for high-quality responses is vital in natural language generation. RLHF (Christiano et al.,, 2017; Ouyang et al.,, 2022) has emerged as a common solution to address this problem by utilizing popular RL methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017a, ). However, this approach faces limitations due to training instability and the complexity introduced by its two-stage pipeline. DPO (Rafailov et al.,, 2023) overcomes this limitation by defining the preference loss as a function of the policy directly given the pairwise preference data. Followup works have extended this algorithm to utilize multiple ranked responses instead of pairwise preference data (Yuan et al., 2023b, ; Liu et al., 2024a, ; Song et al.,, 2024), and avoid dependence on the reference model, effectively merging the instruction tuning phase and preference optimization phase (Hong et al.,, 2024; Meng et al.,, 2024). GPO (Tang et al.,, 2024) unifies several DPO variants (Azar et al.,, 2023; Liu et al., 2024b, ) into a family of algorithms and provides an empirical analysis of the performance trade-off, but leave the theoretical analysis in a heuristic manner. Notably, Wang et al., (2023) also introduces $f$-divergence into the alignment problem, but the divergences are only applied to the regularization term of the RL formulation. Furthermore, all the above works are built on the reparameterization of reward function by policy likelihood, which cannot be guaranteed in practice (Ji et al.,, 2024). In contrast, our method unifies the algorithm based on the theoretical framework of statistical distribution divergence (Csiszár et al.,, 2004; Liese and Vajda,, 2006), which offers a principled to choose the specific algorithm based on divergence characteristics. EXO (Ji et al.,, 2024) proposes learning the policy by exactly optimizing the RLHF objective that minimizes the reverse KL divergence against the optimal policy, which is a special case of our proposed framework. Xiao et al., (2024) leverages demonstration data for alignment using self-imitation learning. In addition, numerous offline preference optimization works propose different training objectives to emphasize various behaviors (Yuan et al., 2023a, ; Zhao et al.,, 2023; Xu et al.,, 2024; Ethayarajh et al.,, 2024).
@@ 3.3 $f$-divergence
Divergence describes the difference between two probability distributions. A general family of divergences is the $f$-divergences (Csiszár et al.,, 2004; Liese and Vajda,, 2006), also known as the Ali-Silvey distances (Ali and Silvey,, 1966). For a convex function $f:\mathbb{R}^{+}\rightarrow\mathbb{R}$ that is lower-semicontinuous, strictly convex around $1$, and satisfies $f(1)=0$, given two distributions $p$ and $q$, the corresponding $f$-divergence for these two distributions is defined as:
@@ 4.2 Generalized Preference Optimization
Pair-wise Preference Data. The most common preference dataset typically consists of pair-wise completions for each prompt, i.e., setting $K=2$. Since there are two completions $y$ for each prompt $x$, we can denote them as “winning" sample $y_{w}$ and “losing" sample $y_{l}$. Then LABEL:eq:loss_fpo_general can be simplified as:
@@ 5.2 Benchmarking $f$-PO
In this subsection, we further demonstrate the strong performance of our $f$-PO on four popular benchmarks through four different settings.
@@
Table 4: Results on LLM Leaderboard v2 with Llama3-Instruct (8B) and Mistral-Base (7B).
IFEval BBH MATH Lvl 5 GPQA MUSR MMLU-PRO Average
SimPO-Llama3-8B-Instruct 65.04 26.71 2.57 5.82 8.15 27.66 22.66
$f$-PO-Llama3-8B-Instruct 67.69 28.75 4.98 4.92 9.08 28.79 24.04
SimPO-Mistral-7B-Base 47.01 22.33 0.60 4.47 8.03 18.91 16.89
$f$-PO-Mistral-7B-Base 48.46 24.16 1.66 6.38 12.73 19.72 18.85
@@
Evaluation metrics. We evaluate our models using four popular open-ended instruction-following benchmarks: AlpacaEval 2 (Dubois et al.,, 2024), MT-Bench (Zheng et al.,, 2023), Arena-Hard 0.1 (Li et al.,, 2024), and Open LLM Leaderboard v2 (Fourrier et al.,, 2024). These benchmarks assess diverse conversational abilities across various queries. For AlpacaEval 2, we report both the raw win rates (WR) and length-controlled win rates (LC). MT-Bench scores are averaged using GPT-4 and GPT-4-Preview-1106 as judges. Arena-Hard results are reported as win rates against the baseline model.
@@
Results. Table 3 shows the primary evaluation outcomes of Mistral-7B and Llama-3-8B in both Base and Instruct configurations. While all methods effectively boost the performance over the SFT model, our proposed method, $f$-PO, achieves the highest score on 16 out of 20 metrics while being competitive on the rest, which highlights the efficacy and robustness of our technique. Quantitatively, $f$-PO achieves an average improvement of 1.4% over the strongest baseline SimPO on the AlpacaEval 2 length-controlled win rate metric. Interestingly, in the Llama-3-8B Base setup, we are able to obtain a remarkable enhancement of 4.8% (28.2% against 23.4%) in win rate compared with SimPO on Arena-Hard, further demonstrating the superiority of $f$-PO. The results consistently advocate $f$-PO as an effective and broadly applicable LLM alignment approach across a wide suite of models and tasks. The results in Table 4 further verifies the superiority of our $f$-PO, where $f$-PO outperforms SimPO consistently with both models (Llama3-8B-Instruct and Mistral-7B-Base) by a significant margin, leading to  8% relative improvement in average score. Notably, $f$-PO enhances the score of Mistral-7B-Base on the MUSR benchmark from 8.03 to 12.73, and Llama3-8B-Instruct on the IFEval benchmark from 22.66 to 24.04. The results further strengthen our findings of $f$-PO as a general and effective preference optimization approach.
@@ 6 Conclusion
Acknowledgment
We gratefully acknowledge the support of ARO (W911NF-21-1-0125), ONR (N00014-23-1-2159), and Chan Zuckerberg Biohub. We also gratefully acknowledge the support of NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM), IIS-2403318 (III); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB. Minkai Xu thanks the generous support of Sequoia Capital Stanford Graduate Fellowship.
@@ Acknowledgment
$f$-PO: Generalizing Preference Optimization with
$f$-divergence Minimization
@@ A.2 Proof of Theorem 2
[…]
@@ Appendix C Prompt Details
In this section, we provide the specific prompts used in our model evaluation, as described in Sections 5.1 and 5.2.
@@ C.1 Dialogue Generation Evaluation Prompt
SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful. Your response should use the format: Comparison: <one-sentence comparison and explanation>
More helpful: <"A" or "B">
@@ C.2 Summarization Evaluation Prompt
For assessing the quality of summaries, we employ the concise prompt structure from Rafailov et al., (2023):
@@
SECOND, on a new line, state only "A" or "B" to indicate your choice. Your response should use the format: Comparison: <one-sentence comparison and explanation>
Preferred: <"A" or "B">